Comparative financial statements for Weller Corporation, a merchandising company, for the year ending December 31 appear below. The company did not issue any new common stock during the year. A total of 810,000 shares of common stock were outstanding. The interest rate on the bond payable was 12%, the income tax rate was 40%, and the dividend per share of common stock was $0.75 last year and $0.40 this year. The market value of the company’s common stock at the end of the year was $18.50. All of the company’s sales are on account.

Weller Corporation
Comparative Balance Sheet
(dollars in thousands)
This Year Last Year
Assets
Current assets:
Cash $ 2,546 $ 2,770
Accounts receivable, net 12,400 9,150
Inventory 9,750 8,240
Prepaid expenses 1,810 2,120
Total current assets 26,506 22,280
Property and equipment:
Land 6,100 6,100
Buildings and equipment, net 19,300 19,100
Total property and equipment 25,400 25,200
Total assets $ 51,906 $ 47,480
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 9,600 $ 8,350
Accrued liabilities 620 750
Notes payable, short term 310 310
Total current liabilities 10,530 9,410
Long-term liabilities:
Bonds payable 6,250 6,250
Total liabilities 16,780 15,660
Stockholders' equity:
Common stock 810 810
Additional paid-in capital 4,250 4,250
Total paid-in capital 5,060 5,060
Retained earnings 30,066 26,760
Total stockholders' equity 35,126 31,820
Total liabilities and stockholders' equity $ 51,906 $ 47,480
Weller Corporation
Comparative Income Statement and Reconciliation
(dollars in thousands)
This Year Last Year
Sales $ 80,000 $ 75,000
Cost of goods sold 52,500 48,500
Gross margin 27,500 26,500
Selling and administrative expenses:
Selling expenses 8,600 8,100
Administrative expenses 12,100 11,100
Total selling and administrative expenses 20,700 19,200
Net operating income 6,800 7,300
Interest expense 750 750
Net income before taxes 6,050 6,550
Income taxes 2,420 2,620
Net income 3,630 3,930
Dividends to common stockholders 324 648
Net income added to retained earnings 3,306 3,282
Beginning retained earnings 26,760 23,478
Ending retained earnings $ 30,066 $ 26,760

Required:

Compute the following financial data for this year:

1. Earnings per share. (Round your answer to 2 decimal places.)

2. Price-earnings ratio. (Round your intermediate calculations and final answer to 2 decimal places.)

3. Dividend payout ratio. (Round your intermediate calculations and final answer to 2 decimal places.)

4. Dividend yield ratio. (Round your answer to 2 decimal places.)

5. Book value per share. (Round your answer to 2 decimal places.)
Answer #1

1. Earnings per share = Net income - preferred stock dividend / Number of equity shares outstanding

Net income for this year = $ 3630000

Number of equity shares outstanding = 810000

therefore Earnings per share = 3630000/ 810000 = $ 4.48 per share

2. Price - earnings ratio = Market price per share / Earnings per share

Market value of company's common stock at the end of the year = $ 18.5

Earnings per share = $ 4.48( calculated on previous question)

Therefore price - earnings ratio = 18.5 / 4.48 = 4.13

3. Dividend payout ratio = (Dividend per share / Earnings per share) * 100

Dividend per share = $.40 per share

earnings per share = $ 4.48 ( calculated as per Question 1)

Therefore Dividend payout ratio = (.40 / 4.48 ) * 100 = 8.93%

4. Dividend yield ratio = (Dividend per share / market value per share) * 100

Dividend per share = $ .40 per share

Market value per Share = $ 18.5 per share

Therefore Dividend yield ratio = (.40 / 18.5) * 100 = 2.16%

5. Book value per share = Total stock holders equity / Number of shares of common stock outstanding

Total stock holders equity = $ 35126000 (given in the question itself)

Number of common stock outstanding = 810000

Book value per share = 35126000 / 810000 = $ 43.37